OpenStudy (anonymous):

please show steps in the simplest form so I can understand. x+7/6=2x-7/6

OpenStudy (anonymous):

combine likes terms x-2x=-7/6-7/6 so -x=-14/12 so -x=-7/6 you ans is x=7/6

OpenStudy (anonymous):

7/6 are in fraction form

OpenStudy (anonymous):

\[x=\frac{7}{3}\] Move the x to the right side and the -7/6 to the left side and simplify. 7/6 + 7/6 = x x= 14/6 or 7/3

OpenStudy (anonymous):

thanks you made it so easy i do understand,